# Nav-SCOPE: Swarm Robot Cooperative Perception and Coordinated Navigation

### v2

@@ VI Real World Experiments
Specifically, we observe convergence effects at points $A$, $B$ and $C$, and a divergence effect at point $D$. All of these are accomplished by incorporating supplementary observation from cooperative perception. A particular case at point $A$ is presented in Figure 8, where it is difficult for $R_{3}$ to decide its future direction, as the current observation looks similar on either side. But after fusion, it chose to follow $R_{4}$ rightwards, and there proved to be fewer obstacles. The swarm also gains advantages from the collision avoidance effect by eliminating path interferences which evidently occurred at points $D$ in the Nav-I baseline. In contrast, Nav-SCOPE shows organized trajectories despite robot interactions and narrow pathways.
VII Conclusion
In this paper, we proposed a systematic solution to achieve cooperative perception and coordinated navigation on swarm robots. Our approach facilitates information flow for observation fusion, followed by path optimization with self-organized convergence, divergence, and collision avoidance. Baseline comparisons demonstrate its exceptional performance, robustness, and lightweight demands. Future work will aim to enhance the smoothness and continuity of trajectories and to test the model on aerial robots such as quadrotors.

### v3

@@ VII Real World Experiments
Specifically, we observe convergence effects at points $A$, $B$ and $C$, and a divergence effect at point $D$. All of these are accomplished by incorporating supplementary information from cooperative perception. A particular case at point $A$ is presented in Figure 9, where it is difficult for $R_{3}$ to decide its future direction, as the current observation looks similar on either side. But after fusion, it chose to follow $R_{4}$ rightwards, and there proved to be fewer obstacles. The swarm also gains advantages from collision avoidance by eliminating path interferences which evidently occurred at points $D$ in the Nav-I baseline. In contrast, Nav-SCOPE shows organized trajectories despite robot interactions and narrow pathways.
VIII Discussion and Conclusion
Despite the current success, we acknowledge that Nav-SCOPE is only applicable to ground robots at this stage, while extension is considered for applications in 3D space. Our future aim is to implement Nav-SCOPE on aerial robots, such as quadrotors. In the perception stage, we will introduce 2D FFT filtering to process depth images [33]. In the planning stage, we will extend the probability function from bi-categories to multi-categories, and incorporate trajectory optimization with robot dynamic costs [34]. In execution, we will add the time-delay calibration in high-speed scenarios [20].
Other promising areas include model adaptation to dynamic environments and task expansion to exploration applications. Some existing theories such as obstacle tracking [35] and target allocation [36], etc. could be incorporated to improve system performance in specific tasks.
In summary, we propose a systematic solution to achieve cooperative perception and coordinated navigation on swarm robots. Our approach facilitates information flow for observation fusion, followed by path optimization with self-organized convergence, divergence, and collision avoidance. Baseline comparisons demonstrate its exceptional performance, robustness, and lightweight demands.